Be sure to answer all parts.

Calculate the pH of each of the following solutions.

(a) 6.9

×

10−4M Ba(OH)2:

×

10

(Enter your answer in scientific notation.)

(b) 2.5

×

10−4M HNO3:
Answer #1

a)

[OH-] = 2*[Ba(OH)2]

= 2*6.9*10^-4 M

= 1.38*10^-3 M

use:

pOH = -log [OH-]

= -log (1.38*10^-3)

= 2.86

use:

PH = 14 - pOH

= 14 - 2.86

= 11.14

Answer: 11.14

b)

[H+] = 2.5*10^-4 M

use:

pH = -log [H+]

= -log (2.5*10^-4)

= 3.60

Answer: 3.60